OpenStudy (anonymous):

solve. 3/4x=18 please show work

OpenStudy (anonymous):

divide 18 by 3/4... aka multiply it by 4/3. 72/3 x=24

OpenStudy (anonymous):

3/4x=18 18*4/3=x x=24

OpenStudy (anonymous):